# Meta Pruning via Graph Metanetworks : A Universal Meta Learning Framework for Network Pruning

### v2

@@ 1 Introduction
With the rapid advancement of deep learning (LeCun et al., 2015; Schmidhuber, 2015), neural networks have become increasingly powerful. However, this improved performance often comes with a significant increase in the number of parameters and computational cost (FLOPs). As a result, there is growing interest in methods to simplify these networks while preserving their performance. Pruning, which involves selectively removing certain parts of a neural network, has proven to be an effective approach (Cheng et al., 2024; He & Xiao, 2024; Reed, 1993).
A key idea of a large number of previous pruning works is to remove the unimportant components of a neural network. To achieve this goal, various criteria that measure the importance of components of a neural network have been invented. Some criteria are based on norm or magnitude scores (Han et al., 2016; Li et al., 2017b; Sun et al., 2024; He et al., 2018a). Some are based on how much the performance changes when certain parts of the network are removed or preserved (You et al., 2019b; Ma et al., 2023; Frankle & Carbin, 2019; Ye et al., 2020). Some others defined more complex criteria such as saliency and sensitivity (LeCun et al., 1989; Lee et al., 2019; Zhao et al., 2019).
@@
Prior researchers have generally approached pruning improvement in two ways: either by refining the pruning criteria or by transforming the model to be easier to prune according to a given criterion. The first approach is more straightforward and has been extensively studied in prior work. In contrast, the second approach has received comparatively less attention. However, several studies adopting this latter perspective have achieved promising results. Notably, sparsity regularization methods (Wen et al., 2016; Fang et al., 2023; Wang et al., 2021) exemplify this approach by encouraging models to become sparser, thereby facilitating more effective pruning.
@@ 2.1 Metanetworks
Using neural networks to process other neural networks has emerged as an intriguing research direction. Early studies demonstrated that neural networks can extract useful information directly from the weights of other networks (Unterthiner et al., 2020). We define a metanetwork as a neural network that takes another neural network as input and outputs either information about it or a modified network. Initial metanetworks were simple multilayer perceptrons (MLPs), while more recent designs incorporate stronger inductive biases by preserving symmetries (Godfrey et al., 2022) inherent in the input networks. Broadly, metanetwork architectures can be categorized into two perspectives. The weight space view applies specially designed MLPs directly on the model’s weights (Zaheer et al., 2017; Navon et al., 2023; Zhou et al., 2023a; Tran et al., 2024b; Zhou et al., 2023b; 2024; Tran et al., 2024a). And the graph view transforms the input network into a graph and applies graph neural networks (GNNs) to it (Lim et al., 2024; Kofinas et al., 2024; Kalogeropoulos et al., 2024). In this work, we adopt the graph view to build our metanetwork, leveraging the natural correspondence between graph nodes and neurons in a network layer, as well as the symmetries they exhibit (Maron et al., 2019). Given the maturity and effectiveness of GNNs, the graph metanetwork approach offers both elegance and practicality.
@@ 2.2 Graph Neural Networks
Graph neural networks (GNNs) are a class of neural networks specifically designed to operate on graph-structured data, where graphs consist of nodes and edges with associated features (Wu et al., 2021; Scarselli et al., 2009; Kipf & Welling, 2017). GNNs have gained significant attention in recent years, with frameworks such as the message passing neural network (MPNN) (Gilmer et al., 2017) proving to be highly effective. In this paper, we employ the message passing framework, specifically using Principal Neighborhood Aggregation (PNA) (Corso et al., 2020) as the backbone architecture for our metanetwork.
@@ 2.3 Sparsity Regularization Based Pruning
Networks that are inherently sparse tend to be easier to prune effectively. Prior research has explored various regularization techniques to encourage sparsity in neural networks to facilitate pruning. This idea of sparsity regularization has been widely adopted across many works (Wen et al., 2016; Gordon et al., 2018; Huang & Wang, 2018; Lin et al., 2020c; He et al., 2017; Lin et al., 2019; Xia et al., 2022; 2024; Fang et al., 2023; Wang et al., 2021). Our method inherits this principle to some extent, as our metanetwork can be viewed as transforming the original network into a sparser version for better pruning.
@@ 2.4 Learning to Prune & Meta Pruning
Given the complexity of pruning, several previous works have leveraged neural networks (Dery et al., 2024; He et al., 2017; Wu et al., 2024; Chen et al., 2023), reinforcement learning (Rao et al., 2019; Yu et al., 2022; He et al., 2018b), and other techniques to automatically learn pruning strategies. Viewing pruning as a learning problem naturally leads to the idea of meta-learning, which learns “how to learn” (Hospedales et al., 2022). Some prior works have specifically applied meta-learning to pruning (Liu et al., 2019; Li et al., 2020). Our work also draws idea from meta-learning, but it learns in a way entirely different from all prior works.
@@ 4.1 Metanetwork Design
Conversion between networks and graphs: To apply a metanetwork (GNN) on networks, we first need to establish a conversion between networks and graphs. Prior works such as Kofinas et al. (2024); Lim et al. (2024) have managed to change networks of almost all architectures (CNN, Transformer, RNN, etc.) into graphs.
Here we show how we establish conversion between a ResNet (He et al., 2016) and a graph in our experiments as an example. This inherits the idea of Kofinas et al. (2024); Lim et al. (2024) but is quite different in implementation details. To convert a network into a graph, we establish the correspondence between the components of a neural network and elements of a graph as follows:
(1) Node: Each neuron in fully connected layers or channel in convolutional layers in the network is represented as a node in the graph. (2) Edge: An edge exists between two nodes if there is a direct connection between them in the original network. This includes fully connected layers, convolutional layers, and residual connections. (3) Node Features: Node features comprise parameters associated with neurons, here we use the weight, bias, running mean, running var of batchnorm (Ioffe & Szegedy, 2015), including both batchnorms from previous adjacent layers and previous skip connection layers (if exist). (4) Edge Features: Edge features encode the weights of connections between nodes. We treat linear connections and residual connections as special cases of convolutional connections with a kernel size of 1. For a $k\times k$ convolution between two channels, we flatten it into a $k^{2}$-dim edge feature vector.
@@
Metanetwork architecture: After establishing the conversion between networks and graphs, we can convert the input network into a graph, use metanetwork(GNN) to transform the graph, and finally convert the output graph back into a new network. Our metanetwork must be powerful enough to learn the transforming rules for better pruning. We build our metanetwork based on the message passing framework (Gilmer et al., 2017) and PNA (Corso et al., 2020) architecture. A brief overview of our architecture is provided below, the full description is available in Appendix B.1.
@@ 4.3 Pruning Criterion
We draw ideas from prior sparsity regularization based pruning methods(section 2.3), and design our pruning criterion based on them (especially Fang et al. (2023)). We will give a brief introduction below. For a rigorous mathematical definition and more details, please refer to appendix B.2.
We only do structural pruning in all our experiments. A variant of the classical $\ell_{2}$ norm, the group norm is used as our importance score function. It inherits the idea of using $\ell_{2}$ norm to calculate a score, but calculate it on a pruning group level. Since structural pruning is performed on groups, traditional $l_{2}$ norm that treats nodes and edges in isolation is no longer viable. Group $l_{2}$ norm that takes a whole group into consideration has shown to be more effective. We calculate the sparisty loss by multiply our group norm scores with some coefficients.
The choice of pruning criteria is highly flexible (see in our later experiments, section 6.1), which further demonstrates the generality of our framework. Although we conduct only structural pruning in all our experiments, all our methods can be directly applied to unstructured pruning as well.
@@ 4.4 A Holistic Perspective
Universally applicability: Theoretically, our implementations can be applied to almost all types of networks with all kinds of pruning. Almost all types of networks can be converted into graphs (Kofinas et al., 2024; Lim et al., 2024), and the designs of the GNN and the pruning criteria are also broadly compatible. This makes our framework a universally applicable solution for network pruning.
@@ 5.1 Preliminaries
Unlike prior learning to prune methods that require special training during each pruning, our pruning pipeline only requires a feed forward through metanetwork and standard finetuning, which is simple, general and effective. Once we have a metanetwork, we can prune as many networks as we want. All pruning models are completely different from meta-training data models. This demonstrates our metanetwork naturally has great transferability and no prior work has done something like this before. See Appendix A for more comparisons between our work and prior ones.
@@ 5.2 Classical CNN Pruning Tasks
We carry out our experiments on three most classical, popular, and representative image recognition tasks, including pruning ResNet56 (He et al., 2016) on CIFAR10 (Krizhevsky & Hinton, 2009), VGG19 (Simonyan & Zisserman, 2015) on CIFAR100 (Krizhevsky & Hinton, 2009) and ResNet50 (He et al., 2016) on ImageNet (Deng et al., 2009). See appendix D.1 for more general setups.
@@
5.3 Ablation Study
@@ 6.1 Flexible Pruning Criterion
Pruning criterion of our method is highly flexible. We build a series of reasonable pruning criterion (Appendix B.2). At first, we directly use MEAN REDUCE, MAX NORMALIZE, $\alpha=4$ as default, because this is the same as group norm in Fang et al. (2023). Later, we keep everything else the same, and try many different criteria, and find they all works well (Table 2). This demonstrates that our framework is robust and has many flexibilities.
@@ 6.2 Transfer between Datasets and Architectures
Section 4.4 mentioned that our metanetwork has natural transferability. Following experiments demonstrate that our metanetwork can transfer between similar datasets and network architectures.
Transfer between datasets: 3 datasets, CIFAR10 (Krizhevsky & Hinton, 2009), CIFAR100 (Krizhevsky & Hinton, 2009) and SVHN (Goodfellow et al., 2014) are used. We train metanetwork and the to be pruned network on each of them, traverse through every possible combination, and use the metanetwork to prune the network. Results (Table 3) show that out metanetwork can transfer between similar datasets. See Appendix F.1 for full details
@@
Transfer between architectures: 2 architectures, ResNet56 and ResNet110 (He et al., 2016) are used. We train metanetwork and the to be pruned network on each of them, traverse through every possible combination, and use the metanetwork to prune the network. Results (Table 4) show that our metanetwork can transfer between similar datasets. See Appendix F.2 for full details
@@ 7 Experiments on Transformers
As mentioned in section 4.4, our framework is theoretically applicable to almost all types of networks with all kinds of pruning. Here we expand our implementations to another widely used arthitecture–transformer (Vaswani et al., 2017). More specifically, we pruned the vision transformer (ViT) from Dosovitskiy et al. (2021) that is trained on ImageNet. See Appendix G for full experiments, including how we convert transformer into graph and conduct further meta-training and pruning.
@@ A.2 Computational and Memory Costs
One important question for almost all previous meta-learning approaches is the computational and memory costs are too large. We provide some estimate of memory and computational costs for scaling our methods to larger models and compare with two classic pruning methods–pruning at initalization methods (cheap pruning, hard to training and low accuracy like Lee et al. (2019); Wang et al. (2022)) and iterative magnitude pruning (costly, pruning during training and high performance like Frankle & Carbin (2019); Molchanov et al. (2019)). The results show that our method achieves outstanding results while uses relatively low costs.
@@ A.3 Generality
Almost all previous meta-learning based pruning approaches, such as Liu et al. (2019); Li et al. (2020); Wu et al. (2024), are tailored to a specific network and thus lack generality. Because of this they require specialized training for each pruning instance.

### v3

@@ 1 Introduction
With the rapid advancement of deep learning (DeepLearning; DeepLearningOverview), neural networks have become increasingly powerful. However, this improved performance often comes with a significant increase in the number of parameters and computational cost (FLOPs). As a result, there is growing interest in methods to simplify these networks while preserving their performance. Pruning, which involves selectively removing certain parts of a neural network, has proven to be an effective approach (PruningSurvey; PruningSurveyStructured; PruningSurveyOld).
A key idea of a large number of previous pruning works is to remove the unimportant components of a neural network. To achieve this goal, various criteria that measure the importance of components of a neural network have been invented. Some criteria are based on norm or magnitude scores (DeepCompression; PruningFiltersForEfficientConvNets; NormLLM; SFP). Some are based on how much the performance changes when certain parts of the network are removed or preserved (GateDecorator; LLMPruner; LotteryTicket; SubnetworkPruning). Some others defined more complex criteria such as saliency and sensitivity (OptimalBrainDemage; Snip; VariationalConvolutionalNeuralNetworkPruning).
@@
Prior researchers have generally approached pruning improvement in two ways: either by refining the pruning criteria or by transforming the model to be easier to prune according to a given criterion. The first approach is more straightforward and has been extensively studied in prior work. In contrast, the second approach has received comparatively less attention. However, several studies adopting this latter perspective have achieved promising results. Notably, sparsity regularization methods (LearningStructuredSparsityInDeepNeuralNetworks; DepGraph; Greg) exemplify this approach by encouraging models to become sparser, thereby facilitating more effective pruning.
@@ 2.1 Metanetworks
Using neural networks to process other neural networks has emerged as an intriguing research direction. Early studies demonstrated that neural networks can extract useful information directly from the weights of other networks (AccFromWeights). We define a metanetwork as a neural network that takes another neural network as input and outputs either information about it or a modified network. Initial metanetworks were simple multilayer perceptrons (MLPs), while more recent designs incorporate stronger inductive biases by preserving symmetries (SymmetryDeepLearning) inherent in the input networks. Broadly, metanetwork architectures can be categorized into two perspectives. The weight space view applies specially designed MLPs directly on the model’s weights (DeepSets; DWS; NFN; NFNTransformer; NFTransformer; UNF; MonomialNFN). And the graph view transforms the input network into a graph and applies graph neural networks (GNNs) to it (GMN; NGGNN; ScaleGMN). In this work, we adopt the graph view to build our metanetwork, leveraging the natural correspondence between graph nodes and neurons in a network layer, as well as the symmetries they exhibit (InvariantEquivariantGNN). Given the maturity and effectiveness of GNNs, the graph metanetwork approach offers both elegance and practicality.
@@ 2.2 Graph Neural Networks
Graph neural networks (GNNs) are a class of neural networks specifically designed to operate on graph-structured data, where graphs consist of nodes and edges with associated features (GNNSurvey; TheGraphNeuralNetworkModel; SemiSupervisedClassificationWithGraphConvolutionalNetworks). GNNs have gained significant attention in recent years, with frameworks such as the message passing neural network (MPNN) (MessagePassing) proving to be highly effective. In this paper, we employ the message passing framework, specifically using Principal Neighborhood Aggregation (PNA) (PNA) as the backbone architecture for our metanetwork.
@@ 2.3 Sparsity Regularization Based Pruning
Networks that are inherently sparse tend to be easier to prune effectively. Prior research has explored various regularization techniques to encourage sparsity in neural networks to facilitate pruning. This idea of sparsity regularization has been widely adopted across many works (LearningStructuredSparsityInDeepNeuralNetworks; MorphNet; DataDrivenSparseStructureSelection; TowardCompactConvNetsViaStructureSparsityRegularizedFilterPruning; ChannelPruningForAcceleratingVeryDeepNeuralNetworks; TowardsOptimalStructuredCNNPruningViaGenerativeAdversarial; StructuredPruningLearnsCompactAndAccurateModels; ShearedLLaMA; DepGraph; Greg). Our method inherits this principle to some extent, as our metanetwork can be viewed as transforming the original network into a sparser version for better pruning.
@@ 2.4 Learning to Prune & Meta Pruning
Given the complexity of pruning, several previous works have leveraged neural networks (EverybodyPruneNow; ChannelPruningForAcceleratingVeryDeepNeuralNetworks; ATO; OTO), reinforcement learning (RuntimeNetworkRoutingForEfficientImageClassification; GNN-RL; AMC), and other techniques to automatically learn pruning strategies. Viewing pruning as a learning problem naturally leads to the idea of meta-learning, which learns “how to learn” (MetaLearningSurvey). Some prior works have specifically applied meta-learning to pruning (MetaPruningFormer; DHP). Our work also draws idea from meta-learning, but it learns in a way entirely different from all prior works.
@@ 4.1 Metanetwork Design
Conversion between networks and graphs: To apply a metanetwork (GNN) on networks, we first need to establish a conversion between networks and graphs. Prior works such as NGGNN; GMN have managed to change networks of almost all architectures (CNN, Transformer, RNN, etc.) into graphs.
Here we show how we establish conversion between a ResNet (Resnet) and a graph in our experiments as an example. This inherits the idea of NGGNN; GMN but is quite different in implementation details. To convert a network into a graph, we establish the correspondence between the components of a neural network and elements of a graph as follows:
(1) Node: Each neuron in fully connected layers or channel in convolutional layers in the network is represented as a node in the graph. (2) Edge: An edge exists between two nodes if there is a direct connection between them in the original network. This includes fully connected layers, convolutional layers, and residual connections. (3) Node Features: Node features comprise parameters associated with neurons, here we use the weight, bias, running mean, running var of batchnorm (BatchNorm), including both batchnorms from previous adjacent layers and previous skip connection layers (if exist). (4) Edge Features: Edge features encode the weights of connections between nodes. We treat linear connections and residual connections as special cases of convolutional connections with a kernel size of 1. For a $k\times k$ convolution between two channels, we flatten it into a $k^{2}$-dim edge feature vector.
@@
Metanetwork architecture: After establishing the conversion between networks and graphs, we can convert the input network into a graph, use metanetwork(GNN) to transform the graph, and finally convert the output graph back into a new network. Our metanetwork must be powerful enough to learn the transforming rules for better pruning. We build our metanetwork based on the message passing framework (MessagePassing) and PNA (PNA) architecture. A brief overview of our architecture is provided below, the full description is available in Appendix B.1.
@@ 4.3 Pruning Criterion
We perform most of our experiments by default using structural pruning, but also include a small number of experiments with unstructured and N:M sparsity pruning to demonstrate that our method can be applied to all kinds of pruning. In unstructured pruning and N:M sparisity pruning, we simply use the naive $l_{1}$ norm as both our importance score function and sparsity loss.
In structural pruning, we draw ideas from prior sparsity regularization based pruning methods(section 2.3), and design our pruning criterion based on them (especially DepGraph). We will give a brief introduction below. For a rigorous mathematical definition and more details, please refer to appendix B.2.
A variant of the classical $\ell_{2}$ norm, the group norm is used as our importance score function. It inherits the idea of using $\ell_{2}$ norm to calculate a score, but calculate it on a pruning group level. Since structural pruning is performed on groups, traditional $l_{2}$ norm that treats nodes and edges in isolation is no longer viable. Group $l_{2}$ norm that takes a whole group into consideration has shown to be more effective. We calculate the sparisty loss by multiply our group norm scores with some coefficients.
The choice of pruning criteria is highly flexible (see in our later experiments, section 6.1), which further demonstrates the generality of our framework. We perform most of our experiments by default using structural pruning, but also include a small number of experiments with unstructured and N:M sparsity pruning to demonstrate that our method can be applied to all kinds of pruning.
@@ 4.4 A Holistic Perspective
Universally applicability: Theoretically, our implementations can be applied to almost all types of networks with all kinds of pruning. Almost all types of networks can be converted into graphs (NGGNN; GMN), and the designs of the GNN and the pruning criteria are also broadly compatible. This makes our framework a universally applicable solution for network pruning.
@@ 5.1 Preliminaries
Unlike prior learning to prune methods that require special training during each pruning, our pruning pipeline only requires a feed forward through metanetwork and standard finetuning, which is simple, general and effective. Once we have a metanetwork, we can prune as many networks as we want. All pruning models are completely different from meta-training data models. This demonstrates our metanetwork naturally has great transferability and no prior work has done something like this before. See Appendix A for more comparisons between our work and prior ones (ideas, costs etc.).
In all experiments, we adopt structured pruning by default. Unstructured pruning will be explicitly stated when used (Section 6.2).
@@ 5.2 Classical CNN Pruning Tasks
We carry out our experiments on three most classical, popular, and representative image recognition tasks, including pruning ResNet56 (Resnet) on CIFAR10 (CIFAR), VGG19 (VGG) on CIFAR100 (CIFAR) and ResNet50 (Resnet) on ImageNet (IMAGENET). See appendix D.1 for more general setups.
@@
5.3 Ablation Study of General Behavioral Tendencies
@@
5.4 Ablation Study of Statistics
We compare the statistics between the origin network and the network after metanetwork(finetuning) to find out how metanetwork transforms the network to make it easier to prune. We mainly visualize the $l2$ norm and taylor sensitivity distribution of each layer in Figure 7. Taylor sensitivity here is defined as $w\cdot\Delta L$, it estimates how much the loss will increase if we mask this weight to zero. The larger taylor sensitivity, the more important the weight is and we are unlikely to prune it. More visualization of other statistics are in Appendix H
Figure 7: Statistics: We compare the statistics of the origin network (orange) and the network after metanetwork and finetuning (blue) when pruning ResNet56 on CIFAR10. The first row is layerwise mean $l_{2}$ norm and taylor sensitivity. Then we randomly select several layers and visualize their $l_{2}$ norm distribution in row 2 and taylor sensitivity distribution in row 3. Observation: Row 1 shows both norm and sensitivity drops on average, especially in the latter part of network where contains a large amount of redundancy and can be substantially pruned. Row 2 shows $l_{2}$ norm distribution has been greatly changed. Large norms still exist but more norms tend to be very small. Row 3 shows more parameters become less important under taylor sensitivity.
@@ 6.1 Flexible Pruning Criterion
Pruning criterion of our method is highly flexible. We build a series of reasonable pruning criterion (Appendix B.2). At first, we directly use MEAN REDUCE, MAX NORMALIZE, $\alpha=4$ as default, because this is the same as group norm in DepGraph. Later, we keep everything else the same, and try many different criteria, and find they all works well (Table 2). This demonstrates that our framework is robust and has many flexibilities.
@@
6.2 Unstructured Pruning & N:M Sparsity Pruning
Most of our experiments are conducted using structured pruning as default. Here we conduct unstructured pruning and N:M sparsity pruning to demonstrate that our methods can be use in all kinds of pruning. We don’t use FLOPs to measure the pruning results because unstructured pruning reduces no FLOPs and needs specialized algorithm to accelerate, so we measure the number of parameters instead. See results in Table 3
Table 3: Unstructured & N:M sparsity pruning: Pruning ResNet56 on CIFAR10, meta-train and prune directly with the classical $l_{1}$ norm. From the results we can see (1) our framework also works great on unstructured pruning. (2) Unstructured pruning outperforms structured pruning because it is more flexible.
Methods Pruned Acc Left Params Methods Pruned Acc Left Params
Unstructured 93.95% 50.25% Structured(2.9x) 93.49% 42,96%
94.14% 20.40% 3:4 94.02% 75.14%
93.43% 15.42% 2:4 93.97% 50.27%
92.96% 10.45% 1:4 93.13% 25.41%
@@ 6.3 Transfer between Datasets and Architectures
Section 4.4 mentioned that our metanetwork has natural transferability. Following experiments demonstrate that our metanetwork can transfer between similar datasets and network architectures. More relative experiments are in Appendix F.3.
Transfer between datasets: 3 datasets, CIFAR10 (CIFAR), CIFAR100 (CIFAR) and SVHN (SVHN) are used. We train metanetwork and the to be pruned network on each of them, traverse through every possible combination, and use the metanetwork to prune the network. Results (Table 4) show that out metanetwork can transfer between similar datasets. See Appendix F.1 for full details
@@
Transfer between architectures: 2 architectures, ResNet56 and ResNet110 (Resnet) are used. We train metanetwork and the to be pruned network on each of them, traverse through every possible combination, and use the metanetwork to prune the network. Results (Table 5) show that our metanetwork can transfer between similar datasets. See Appendix F.2 for full details
@@ 7 Experiments on Transformers
As mentioned in section 4.4, our framework is theoretically applicable to almost all types of networks with all kinds of pruning. Here we expand our implementations to another widely used arthitecture–transformer (Transformer). More specifically, we pruned the vision transformer (ViT) from ViT that is trained on ImageNet. See Appendix G for full experiments, including how we convert transformer into graph and conduct further meta-training and pruning.
@@ A.2 Computational and Memory Costs
One important question for almost all previous meta-learning approaches is the computational and memory costs are too large. We provide some estimate of memory and computational costs for scaling our methods to larger models and compare with two classic pruning methods–pruning at initalization methods (cheap pruning, hard to training and low accuracy like Snip; pruningatinitsurvey) and iterative magnitude pruning (costly, pruning during training and high performance like LotteryTicket; ImportanceEstimationforNNP). The results show that our method achieves outstanding results while uses relatively low costs.
@@ A.3 Generality
Almost all previous meta-learning based pruning approaches, such as MetaPruningFormer; DHP; ATO, are tailored to a specific network and thus lack generality. Because of this they require specialized training for each pruning instance.
@@
We divided the generality of learning to prune methods into 4 stages.
  1. (1)
    Learning once prune one specific network.
  2. (2)
    Learning once prune one type of networks(same architecture and dataset) as many as we want.
  3. (3)
    Learning once prune one group of networks(similar architectures and datasets) as many as we want.
  4. (4)
    Learning once prune any networks.
Here learning refers to learn with extra trainings like gradient descent, reinforcement learning, etc. Finetuning or other fixed rules processes are not included. As far as we know, all prior learning to prune methods only reach stage (1). Our method reaches stage (3) and shows great improvement in generality over prior works. Follow our ideas and pretrain the metanetwork on various networks and datasets may provide a possible way to state (4), whether this will work requires further exploration in the future.
In summary, our method shows great improvement in generality over prior learning to prune methods.
A.4 A Concrete Example
We provide a concrete example compared with prior works and report everything-time, hardware, gpu memory, results, etc. to give readers a more intuitive understanding of our method.
We compare our work with DepGraph on pruning ResNet56 on CIFAR10. We choose DepGraph because we want to compare our work with learning to prune methods in recent years that also have strong results like us. The best candidates are DepGraph and ATO. While ATO is a pruning before training method, both DepGraph and our work are pruning after training methods, so we choose DepGraph. The key idea of DepGraph is sparsity training before pruning. It does special training on the network before pruning to make it more sparse and easier to prune.
All experiments are run on 1 NVIDIA RTX 4090. See table 7 for time consumptions of dfferent methods. To align with DepGraph, we don’t use init pruning, and target at a speed up of 2.5x, which is the largest speed up in paper DepGraph. In ours(full), we finetune 100 epochs after metanetwork and 100 epochs after pruning. But later we found our method is stronger and the speed up 2.5x isn’t that hard. So in ours(efficient), we finetune 60 epochs after metanetwork and 60 epochs after pruning but also get results comparable to DepGraph. The DepGraph experiments use their default settings, sparsity training 100 epochs before pruning and finetune 100 epochs after pruning.
Table 7: Time (miniutes)
Process Ours(full) Ours(efficient) DepGraph
Meta-Train 192 + 165 = 357 192 + 165 = 357 0
Prune & Finetune 67 43 84
From the results we can see our methods is quite efficient. It uses 357 minutes for meta-train, which is in a resonable range. We can greatly reduce time consumption in this process with some experiences, but for the sake of fairness in comparison, we must pretend we have no relevant experience. Among the 357 minutes, 192 minutes are used for generating 100 epochs metanetworks and 165 minutes are used for visualizing the Acc VS. Speed Up Curve to select the appropriate metanetwork for pruning. Once a metanetwork is trained, it can be used to prune as many networks as we want, so we can amortize the time for meta-training into each pruning. Figure 8 shows the amortized time consumption of our work compared to DepGraph. In all, our work is more powerful and more efficient in time if we prune several more networks or already have a trained metanetwork.
Figure 8: Amortized Time Consumption
As we mentioned in section A.2, our methods use a large VRAM for a long time in meta-training, a still large but relatively smaller VRAM for a very short time when feed forward through metanetwork during pruning, and a small VRAM for the standard finetuning. Specifically in this experiment, we use 10000 MiB for meta-training, 6000 MiB for feedforward through metanetwork, and 1000 MiB for finetuning. NVIDIA A100 with 80 GiB VRAM is enough for meta-training large models like ResNet50 and ViT-B-16. When we don’t have enough memory capacity, we can also change the batch size and use more time to make up for the lack of our memory capacity. We’ve also tried reducing the size of metanetwork and fit on NVIDIA A100 40G and the results seem not much influenced. All finetuning are standard finetuning and can be done on NVIDIA RTX 4090 with 24G VRAM. In all, during pruning only the feed forward through metanetwork requires large VRAM for a very short time and anything else is plain finetuning and requires small VRAM, during meta-training the VRAM is in a resonable range.